# Fundamental physical constants, operation of physical phenomena and entropy increase

### v5

@@ 3.3 Superconducting transition
Another interesting realisation comes from observing that a variation of FPCs giving $T_{c}^{u}$ of, for example, 105 K or higher, would imply that many materials would have $T_{c}$ above room temperature. Consistent with Eq. (18) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to current FPCs, we observe superconductivity in the range $T\lesssim$ 100-200 K, with the highest $T_{c}$ seen in hydrides [49, 50, 51]. This naturally stimulates the current intense research into finding materials with $T_{c}$ of 300 K and above. We therefore see that the very existence of the current research to find systems with $T_{c}$ above 300 K is itself due to the values of FPCs currently observed.
3.4 Transitions between states of matter: solids, liquids and gases
@@
The triple point pressure of many substances is below the atmospheric pressure, and so we often observe melting and boiling transitions on temperature increase. The phase co-existence lines in Figure 3 are defined by the equality of thermodynamic potentials. At high pressure, melting temperature substantially depends on pressure (see Ref. [52] for the functional form of pressure-temperature melting lines). At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy [53] because microscopically, melting involves breaking chemical bonds. This cohesive energy is set by FPCs in Eq. (4). Melting temperatures correlate with cohesive energies and are on the order of $\omega_{m}$ or $\omega_{\rm D}$ in the range 102-103 K in many systems [53, 54] (the difference between the cohesive energy and $\hbar\omega_{m}$ is due to the small factor $\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ in Eq. (11)). As discussed in Section 3.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs. This gives the following role of FPCs in setting the states of matter.
@@
We see that, similarly to structural phase transitions discussed in Section 3.2, the current values of FPCs set the observability of transitions between different states of matter in Figure 3 and the existence of these states.
The apparent overlap of our environmental conditions and melting and boiling temperatures set by FPCs gives rise to two states important for life:
@@
A more detailed consideration shows an additional and more specific role of FPCs in life processes: FPCs have a biofriendly window constrained by biofriendly viscosity and diffusion setting the motion in essential processes in and across cells [32, 33].
We note that classical entropies of solids and gases contain FPCs explicitly. Recalling the solid entropy $S_{s}\propto 3N\ln\frac{T}{\hbar\omega_{\rm D}}$ [46], $\omega_{\rm D}\approx\omega_{m}$ and $\omega_{m}$ in Eq. (12) with $m\propto m_{p}$ gives the dependence of $S_{s}$ on FPCs as
@@
The entropy of the classical gas is [46]
@@
Similarly to what we saw in Section 3.2 for structural phase transitions, the emerging high-temperature gas phase increases the entropy in comparison to what the entropy of the solid would have been in the absence of phase transitions (for first-order transitions, the entropy increase has a contribution from the latent heat). Similarly to structural transitions, transitions between different states of matter are enabled by FPCs as discussed earlier in this section.
@@
Differently from solid and gas entropies in Eqs. (5), (15) and (20), no equation for the liquid entropy exists due to the fundamental no-small parameter problem of the liquid theory (see Ref. [52] for a quick review and Refs. [55, 56] for a more detailed discussion). Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting [55, 56]. However, the energy, the primary property in statistical physics [46], can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy $E$ is set by the energies of its excitations [57]. In liquids, these excitations are gapped phonons and particle jumps [55, 56, 58, 59, 60, 61]. The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}=\frac{dE}{dT}$ (in almost entire range of liquid relaxation time, $E$ and $c_{v}$ are vibrational to a very good approximation [62]). Nevertheless, liquids are a cohesive state, and our discussion of how different states of matter emerge depending on cohesive energies set by FPCs applies to liquids too.
3.5 Metastable states and chemical reactions
@@
In the previous sections, we discussed phase transitions and the existence of different states as a result of these transitions. The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest free energy minimum. Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref. [63] for review). When discussing transitions in these systems, the tacit assumption is the one we mentioned in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
@@
Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the lowest free energy minimum for a given composition. The world of organic compounds relevant for life processes is nearly all metastable [63]. Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) [64].
@@
Transitions between metastable states involve the same atomistic processes as transitions between equilibrium states and energies related to the chemical bond energy set by FPCs. Hence, similarly to equilibrium transitions, the observability and operation of transitions between different metastable states and the existence of numerous metastable phases is set by FPCs. The entropy of new high-temperature metastable states emerging in these transitions is expected to be higher because these states tend to be more symmetric with more open structure, as is the case for equilibrium transitons.
@@
4 Fundamental constants and entropy increase
We consider that understanding the current values of FPCs requires the discussion of their variation. Theories and observations of this variation were previously discussed [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 4, 33]. Anthropic arguments use this variation to rationalise the observed values of FPCs [9, 10, 13]. A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) [43] and guided by a unique design and an exquisite symmetry [65]. FPCs would then be calculable from this theory.
In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases, increase the entropy. We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and the associated increase of the phase space and entropy. We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy. This suggests that the current values of FPCs give rise to entropy increase.
We discussed the entropy increase in condensed matter systems, however this increase applies to different length and energy scales in the Universe on general grounds. Recall Eq. (16) giving the temperature of phase transitions as a function of FPCs in condensed matter. However, the temperature of transitions and reactions depend on FPCs at other energy and length scales as well [6, 13], enabling transitions and reactions at those scales. Similarly to condensed matter systems, we expect the new states emerging as a result of these transitions to have higher entropy. Hence we expect that FPCs promote entropy increase at different levels of structure in the Universe.
@@ 4 Fundamental constants and entropy increase
where $E_{i}$ are energies of subsystems and (b) $d\rho$ is maximised when $E_{i}$ of all subsystems in their own equilibrium states are equal to their most probable equilibrium values $\bar{E_{i}}$, corresponding to full equilibrium [46]. As discussed in Section 3.2, the subsystem where transitions and reactions take place (phase-changing subsystem, PCS) is not closed and not in equilibrium because these transitions are enabled by the heat flow from another heat-supplying subsystem. However, subsystems can be identified which are (a) larger, include both subsystems and are closed and (b) in equilibrium at long post-reaction times [46]. The conventional mechanism of increase of $S$ and $d\rho$ applies to these closed subsystems, and our entropy increase in the PCS enabled by FPCs additionally increases the statistical probability in Eq. (22). As discussed below, identifying collections of subsystems which can be considered as conditionally closed at different levels of the Universe can be of interest in future work.
@@
We note that the environmental parameters can be variable in some processes. For example, nuclear synthesis reactions in a forming star depend on (environmental) pressure and temperature, which, in turn, depend on the mass of collapsing interstellar gas. This variability can affect transitions and reactions and the properties of emergent structures. In some processes such as the above example of nuclear synthesis in stars, the environmental parameters can also change as a result of reactions themselves. Nevertheless, the transitions and reactions are enabled or disabled by FPCs at varying environmental parameters as they are at fixed parameters.
@@
We recall the discussion of long-lived metastable states with higher entropy in Section 3. As recently reviewed [66], metastable states are ubiquitous at different levels of matter organization in the Universe including nuclei and astronomical objects. The energy barriers setting the lifetime of these states vary in a wide range 10-7 to 109 eV per particle, with relaxation times often exceeding the age of the Universe [66]. These metastable states, similarly to those discussed in Section 3, emerge as a result of different transitions and reactions which depend on the values of FPCs [6, 13] and can co-exist at the same environmental parameters.
@@
The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where this increase is analogous to selection in biological evolution due to environmental pressure. Based on this entropy increase and the associated increase of statistical probability (22), we conjecture that entropy increase is a selection principle for FPCs considered to vary in earlier discussions (see, e.g., Refs. [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 4, 33] and references therein). In other words, FPCs tend to give the most probable state. The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, is unknown, although several proposals were discussed [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 33]. The above conjecture is nevertheless interesting to consider because, if correct, it is analogous to introducing a Darwinian concept of a selection principle, the notion that brought new understanding even though the molecular DNA-based mutation mechanism was not known at the time and came about a century later only.
@@
Cyclic or oscillatory models of the Universe (see, e.g., Refs. [67, 35, 68, 69, 70, 71, 72] and Ref. [73] for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle. In these models, the change of FPCs is often related to the smallest size of the Universe and very high energy including the Planck energy scale where unknown physics operates. The size of the oscillating Universe at the inversion point in Ref. [72] is set by the system energy as a parameter in the model and does not need to be small.
@@
We might ask whether the current FPCs give structures in the Universe that correspond to the maximised phase space and entropy or close to it. Adams concludes [13] that the observable Universe is not the most abundant and complex possible: somewhat different values of FPCs would give more complex and abundant structures (e.g. more baryons and stars, more efficient nuclear synthesis and so on). Another interesting observation in Ref. [13] is that our Universe does not lie at the centre of allowed range of values of FPCs. The evolutionary mechanism with the entropy increase as a selection principle has several potential implications for these observations. We first note that this selection applies to FPCs and FPC-enabled structures existing at the time (as is the case in biological evolution) rather than to currently observed structures. Second, the observed structures depend on the magnitude of fluctuations of FPCs: large fluctuations destroy the intricate network of constraints on the values of FPCs needed for different structures to exist [13], resulting in scarcity of structures and smaller phase space. On the other hand, smaller fluctuations maintain the continuity of structures in different cycles. As a result, different structures can emerge along a sequence of cycles, with the entropy of these structures fluctuating along the sequence. Third, it follows that maximising entropy as a function of FPCs is subject to many constraints on their values. This invites an input from the optimisation theory.
@@
An example related to this discussion and involving constraints is the black hole (BH) entropy. If most of the current entropy in the Universe is associated with the entropy of large BHs [74], one might consider small $\hbar$ and $c$ to maximise the BH entropy
@@
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (2), breaking many constraints required for key processes of structure formation and disrupting those processes [6, 13]. This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times [6, 13] (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs). We also note that considering variation of the BH entropy due to different FPCs is unimportant for the evolution of FPCs at earlier stages when BHs may not have existed yet: as mentioned earlier, this evolution applies to structures existing at the time only.
The discussion of FPCs acting to increase the entropy and the conjecture of entropy increase as a selection principle for FPCs need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe [6, 13], identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument applies, clarifying the role of gravity [46] and so on. Given the scale of the problem we face in understanding the origin of fundamental constants [6, 15, 14], making new observations and asking new questions is useful for progress.
5 Discussion and summary
Approaching the challenge to explain fundamental constants started with understanding the role these constants play. This role has historically been limited to high-energy nuclear physics and astrophysics physics and corresponding structures [6, 7, 8, 9, 10, 11, 12, 13, 3]. Condensed matter physics was relatively unexplored in this regard, however more recently it was realised that FPCs set lower and upper bounds on key condensed matter properties. This has enriched condensed matter physics and deepened our understanding of its key phenomena.
Here, we discussed a much wider role played by FPCs in condensed matter physics: at given environmental parameters, FPCs define observability and operation of entire physical effects and phenomena. Our examples included the operation of different phase transitions and emerging new phases, existence of different states of matter, metastable phases, chemical reactions, their products and so on. These results were based on a theory of each physical effect and process rather than on a dimensional analysis as might have been construed. Ref. [31] (pages 479-480) lists important differences between such a theory-based approach and a dimensional analysis.

### v6

@@ III.3 Superconducting transition
Another interesting realisation comes from observing that a variation of FPCs giving $T_{c}^{u}$ of, for example, 105 K or higher, would imply that many materials would have $T_{c}$ above room temperature. Consistent with Eq. (19) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to current FPCs, we observe superconductivity in the range $T\lesssim$ 100-200 K, with the highest $T_{c}$ seen in hydrides Flores-Livas et al. (2020); Duan et al. (2017); Pickard et al. (2019). This naturally stimulates the current intense research into finding materials with $T_{c}$ of 300 K and above. We therefore see that the very existence of the current research to find systems with $T_{c}$ above 300 K is itself due to the values of FPCs currently observed.
III.4 Transitions between states of matter: solids, liquids and gases
@@
The triple point pressure of many substances is below the atmospheric pressure, and so we often observe melting and boiling transitions on temperature increase. The phase co-existence lines in Figure 2 are defined by the equality of thermodynamic potentials. At high pressure, melting temperature substantially depends on pressure (see Ref. Trachenko (2024) for the functional form of pressure-temperature melting lines). At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy Ubbelohde (1978) because microscopically, melting involves breaking chemical bonds. This cohesive energy is set by FPCs in Eq. (2). Melting temperatures correlate with cohesive energies and are on the order of $\omega_{m}$ or $\omega_{\rm D}$ in the range 102-103 K in many systems Ubbelohde (1978); Grimvall and Sjödin (1974) (the difference between the cohesive energy and $\hbar\omega_{m}$ is due to the small factor $\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ in Eq. (12)). As discussed in Section III.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs. This gives the following role of FPCs in setting the states of matter.
@@
We see that, similarly to structural phase transitions discussed in Section III.2, the current values of FPCs set the observability of transitions between different states of matter in Figure 2 and the existence of these states.
The apparent overlap of our environmental conditions with characteristic melting and boiling temperatures set by FPCs gives rise to two states important for life:
@@
A more detailed consideration shows an additional and more specific role of FPCs in life processes: FPCs have a biofriendly window constrained by biofriendly viscosity and diffusion setting the motion in essential processes in and across cells Trachenko (2023b, c).
We note that classical entropies of solids and gases contain FPCs explicitly. Recalling the solid entropy $S_{s}\propto 3N\ln\frac{T}{\hbar\omega_{\rm D}}$ Landau and Lifshitz (1970), $\omega_{\rm D}\approx\omega_{m}$ and $\omega_{m}$ in Eq. (13) with $m\propto m_{p}$ gives the dependence of $S_{s}$ on FPCs as
@@
The entropy of the classical gas is Landau and Lifshitz (1970)
@@
Similarly to what we saw in Section III.2 for structural phase transitions, the emerging high-temperature gas phase increases the entropy in comparison to what the entropy of the solid would have been in the absence of phase transitions (for first-order transitions, the entropy increase has a contribution from the latent heat). Similarly to structural transitions, transitions between different states of matter are constrained FPCs as discussed earlier in this section.
@@
Differently from solid and gas entropies in Eqs. (6), (16) and (21), no equation for the liquid entropy exists due to the fundamental no-small parameter problem of the liquid theory (see Ref. Trachenko (2024) for a quick review and Refs. Trachenko and Brazhkin (2016); Trachenko (2023d) for a more detailed discussion). Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting Trachenko and Brazhkin (2016); Trachenko (2023d). However, the energy, the primary property in statistical physics Landau and Lifshitz (1970), can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy $E$ is set by the energies of its excitations Lifshitz and Pitaevskii (2006). In liquids, these excitations are gapped phonons and particle jumps Trachenko and Brazhkin (2016); Trachenko (2023d); Proctor (2020, 2021); Chen (2022); de With (2024). The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}=\frac{dE}{dT}$ (in almost entire range of liquid relaxation time, $E$ and $c_{v}$ are vibrational to a very good approximation Trachenko and Brazhkin (2013)). Nevertheless, liquids are a cohesive state, and our discussion of how different states of matter emerge depending on cohesive energies set by FPCs applies to liquids too.
III.5 Metastable states and chemical reactions
@@
In the previous sections, we discussed phase transitions and the existence of different states as a result of these transitions. The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest free energy minimum. Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref. Brazhkin (2006) for review). When discussing transitions in these systems, the tacit assumption is the one we mentioned in Section III.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
@@
Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the lowest free energy minimum for a given composition. The world of organic compounds relevant for life processes is nearly all metastable Brazhkin (2006). Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) Cortes et al. .
@@
Transitions between metastable states involve the same atomistic processes as transitions between equilibrium states. The kinetics of transitions between metastable states is governed by the bond energies. These energies are governed by FPCs as discussed in Section III. Hence, similarly to equilibrium transitions, the observability and operation of transitions between different metastable states and the existence of numerous metastable phases is set by FPCs. The entropy of new high-temperature metastable states emerging in these transitions is expected to be higher because these states tend to be more symmetric with more open structure, as is the case for equilibrium transitons.
@@
IV Fundamental constants and entropy increase
We consider that understanding the current values of FPCs requires the discussion of their variation. Theories and observations of this variation were previously discussed Smolin (1992); Wilczynska et al. (2020); Seto et al. (2023); Webb et al. (2011); King et al. (2012); Chiba (2011); Schellekens (2013); Martins (2017); Barrow (2003); Adams (2019); Uzan (2011); Trachenko (2023c). Anthropic arguments use this variation to rationalise the observed values of FPCs Carr (2009); Sloan et al. (2022); Adams (2019). A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) Weinberg (1993) and guided by a unique design and an exquisite symmetry Zee (2007). FPCs would then be calculable from this theory.
In Section III, we discussed different cases where phase transitions and the emergence of different phases due to FPCs increase the entropy. We saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy. This suggests that the current values of FPCs give rise to entropy increase.
We discussed the entropy increase in condensed matter systems, however this increase applies to different length and energy scales in the Universe on general grounds. Recall Eq. (17) giving the temperature of phase transitions as a function of FPCs in condensed matter. However, the temperature of transitions and reactions depend on FPCs at other energy and length scales as well Barrow (2003); Adams (2019), enabling transitions and reactions at those scales. Similarly to condensed matter systems, we expect the new states emerging as a result of these transitions to have higher entropy. Hence we expect that FPCs promote entropy increase at different levels of structure in the Universe.
@@ IV Fundamental constants and entropy increase
where $E_{i}$ are energies of subsystems and (b) $d\rho$ is maximised when $E_{i}$ of all subsystems in their own equilibrium states are equal to their most probable equilibrium values $\bar{E_{i}}$, corresponding to full equilibrium Landau and Lifshitz (1970). As discussed in Section III.2, the subsystem where transitions and reactions take place (phase-changing subsystem, PCS) is not closed and not in equilibrium because these transitions are enabled by the heat flow from another heat-supplying subsystem. However, subsystems can be identified which are (a) larger, include both subsystems and are closed and (b) in equilibrium at long post-reaction times Landau and Lifshitz (1970). The conventional mechanism of increase of $S$ and $d\rho$ applies to these closed subsystems, and our entropy increase in the PCS due to FPCs additionally increases the statistical probability in Eq. (23). As discussed below, identifying collections of subsystems which can be considered as conditionally closed at different levels of the Universe can be of interest in future work.
@@
We note that the environmental parameters can be variable in some processes. For example, nuclear synthesis reactions in a forming star depend on (environmental) pressure and temperature, which, in turn, depend on the mass of collapsing interstellar gas. This variability can affect transitions and reactions and the properties of emergent structures. In some processes such as the above example of nuclear synthesis in stars, the environmental parameters can also change as a result of reactions themselves. Nevertheless, the transitions and reactions are constrained by FPCs at varying environmental parameters as they are at fixed parameters.
@@
We recall the discussion of long-lived metastable states with higher entropy in Section III. As recently reviewed Brazhkin (2024), metastable states are ubiquitous at different levels of matter organization in the Universe including nuclei and astronomical objects. The energy barriers setting the lifetime of these states vary in a wide range 10-7 to 109 eV per particle, with relaxation times often exceeding the age of the Universe Brazhkin (2024). These metastable states, similarly to those discussed in Section III, emerge as a result of different transitions and reactions which depend on the values of FPCs Barrow (2003); Adams (2019) and can co-exist at the same environmental parameters.
@@
The entropy increase due to FPCs can be discussed in terms of an evolution-like mechanism where this increase is analogous to selection in biological evolution due to environmental pressure. Based on this entropy increase and the associated increase of statistical probability (23), we conjecture that total entropy increase is a selection principle for FPCs considered to vary in earlier discussions (see, e.g., Refs. Smolin (1992); Wilczynska et al. (2020); Seto et al. (2023); Webb et al. (2011); King et al. (2012); Chiba (2011); Schellekens (2013); Martins (2017); Barrow (2003); Adams (2019); Uzan (2011); Trachenko (2023c) and references therein). In other words, FPCs tend to give the most probable state. The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, is unknown, although several proposals were discussed Smolin (1992); Wilczynska et al. (2020); Seto et al. (2023); Webb et al. (2011); King et al. (2012); Chiba (2011); Schellekens (2013); Martins (2017); Barrow (2003); Adams (2019); Trachenko (2023c). The above conjecture is nevertheless interesting to consider because, if correct, it is analogous to introducing a Darwinian concept of a selection principle, the notion that brought new understanding even though the molecular DNA-based mutation mechanism was not known at the time and came about a century later only.
@@
Cyclic or oscillatory models of the Universe (see, e.g., Refs. Adams and Laughlin (1997); Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013); Trachenko (2017) and Ref. Kragh (2009) for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle. In these models, the change of FPCs is often related to the smallest size of the Universe and very high energy including the Planck energy scale where unknown physics operates. The size of the oscillating Universe at the inversion point in Ref. Trachenko (2017) is set by the system energy as a parameter in the model and does not need to be small.
@@
We might ask whether the current FPCs give structures in the Universe that correspond to the maximised phase space and entropy or close to it. Adams concludes Adams (2019) that the observable Universe is not the most abundant and complex possible: somewhat different values of FPCs would give more complex and abundant structures (e.g. more baryons and stars, more efficient nuclear synthesis and so on). Another interesting observation in Ref. Adams (2019) is that our Universe does not lie at the centre of allowed range of values of FPCs. The evolutionary mechanism with the entropy increase as a selection principle has several potential implications for these observations. We first note that this selection applies to FPCs and structures appearing due to these FPCs existing at the time (as is the case in biological evolution) rather than to currently observed structures. Second, the observed structures depend on the magnitude of fluctuations of FPCs: large fluctuations destroy the intricate network of constraints on the values of FPCs needed for different structures to exist Adams (2019), resulting in scarcity of structures and smaller phase space. On the other hand, smaller fluctuations maintain the continuity of structures in different cycles. As a result, different structures can emerge along a sequence of cycles, with the entropy of these structures fluctuating along the sequence. Third, it follows that maximising entropy as a function of FPCs is subject to many constraints on their values. Hence understanding the values of FPCs (e.g. calculating these values or their ranges) invites an input from the optimisation theory. This programme would include taking an initial set of FPCs as variables, calculating the entropies of important structures and formations in the Universe and finding FPCs or their ranges by maximising the total entropy while (a) accounting for constraints on FPCs from each structure Barrow (2003); Adams (2019) and (b) remembering that later-appearing structures may rely on the existence of earlier ones. This process prevents, for example, finding FPCs giving larger entropy of earlier structures but smaller entropy of later structures (or no entropy of later structures if they are not formed due to unfavourable FPCs involved in the emergence of earlier structures) or FPCs giving larger entropy of later structures but neglecting earlier structures needed for later structures to emerge.
@@
An example related to this discussion and involving constraints is the black hole (BH) entropy. If most of the current entropy in the Universe is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
@@
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (4), breaking many constraints required for key processes of structure formation and disrupting those processes Barrow (2003); Adams (2019). This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs).
The discussion of FPCs acting to increase the entropy and the conjecture of entropy increase as a selection principle for FPCs need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe Barrow (2003); Adams (2019), identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument applies, clarifying the role of gravity Landau and Lifshitz (1970) and so on. Given the scale of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.
V Discussion and summary
Approaching the challenge to explain fundamental constants started with understanding the role these constants play. This role has historically been limited to high-energy nuclear physics and astrophysics physics and corresponding structures Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003). Condensed matter physics was relatively unexplored in this regard, however more recently it was realised that FPCs set lower and upper bounds on key condensed matter properties. This has enriched condensed matter physics and deepened our understanding of its key phenomena.
Here, we discussed a much wider role played by FPCs in condensed matter physics: at given environmental parameters, FPCs define observability and operation of entire physical effects and phenomena. Our examples included the operation of different phase transitions and emerging new phases, existence of different states of matter, metastable phases, chemical reactions, their products and so on. These results were based on a theory of each physical effect and process rather than on a dimensional analysis as might have been construed. Ref. Trachenko (2023a) (pages 479-480) lists important differences between such a theory-based approach and a dimensional analysis.